OpenStudy (anonymous):

Use trigonometric Identities to solve: 1.sinx/1+sinx - sinx/1-sinx? 2.(sinx+cosx)^2/1+2sinxcosx?

OpenStudy (anonymous):

= (sin(x)-sin^2(x)-sin(x)-sin^2(x))/1-sin^2(x) = -2sin^2(x)/cos^2(x) = -2tan^2(x)

OpenStudy (anonymous):

= sin^2(x)+cos^2(x)+2sin(x)cos(x)/1+2sin(x)cos(x) = 1+2sin(x)cos(x)/1+2sin(x)cos(x) =1

OpenStudy (anonymous):

(sin(x)-cos(x))^2 =[sin(x)-cos(x))(sin(x)-cos(x)] =sin^2(x)-2sin(x)cos(x)+cos^2(x) =1-2sin(x)cos(x)

OpenStudy (anonymous):

Thanks guys! I had the first one right :)